OpenStudy (mony01):

The position of a 50 g oscillating mass is given by x(t)=(2.0cm)cos(10t−π/4), where t is in s. If necessary, round your answers to three significant figures. Determine: the amplitude 2cm the period .628s the spring constant 5N/m the phase constant -.785rad the initial coordinate of the mass 1.41cm the initial velocity 14.1 cm/s maximum speed 20 cm/s what's the total energy?

OpenStudy (fibonaccichick666):

ode or physics?

OpenStudy (mony01):

physics

OpenStudy (fibonaccichick666):

ah energy, physics

OpenStudy (fibonaccichick666):

ok, what eq ya gonna use?

OpenStudy (mony01):

thats what idk

OpenStudy (fibonaccichick666):

oh it just wants total, so you have to think about all the energy at this point

OpenStudy (fibonaccichick666):

ie is it moving, what type of energy is that? is it stopped?

OpenStudy (fibonaccichick666):

then remember conservation if you have anything you have to solve for

OpenStudy (mony01):

is there a formula?

OpenStudy (fibonaccichick666):

formula for conservation yes, for total energy, other than the obvious, not that I know of

OpenStudy (fibonaccichick666):

^that is why I hate physics :)

OpenStudy (fibonaccichick666):

I assume you know how to find spring energy, yes?

OpenStudy (anonymous):

\[F=\frac{ 1 }{ 2 }kx^2\]

OpenStudy (fibonaccichick666):

you are both welcome to take this off of my hands, I am probably not as good at explaining this as either of you could be

OpenStudy (anonymous):

im not that great either

OpenStudy (aum):

The energy of the oscillating mass is made up of potential energy (stored in the energy of the spring) and kinetic energy. When KE is max, PE is zero. When PE is max, KE is zero. You already found max velocity to an earlier question. 1/2 * m * v^2 is the max. KE and at that instant PE will be zero. So that is the total energy.

OpenStudy (mony01):

so the answer is 0

OpenStudy (aum):

\[ \text{Total Energy } = KE_{\text{max}} = \frac 12mv_{\text{max}}^2 = \frac 12 * \frac{50}{1000} * \left( \frac{20}{100} \right)^2 \]

OpenStudy (fibonaccichick666):

You are using ME=ME=ME here,

OpenStudy (fibonaccichick666):

(that is the deeper concept)

OpenStudy (perl):

what does ME stand for?

OpenStudy (aum):

\[ \text{Total Energy } = KE + PE = KE_{\text{max}} + 0 = 0 + PE_{\text{max}} \]Here we use: \[ \text{Total Energy } = KE_{\text{max}} \]

OpenStudy (fibonaccichick666):

Mechanical energy, but that is just the notation that represents conservation of energy from x=0, x=1/2max, x=max

OpenStudy (mony01):

ok so i solved the top one and got .001

OpenStudy (aum):

0.001 Joules or 0.001 N-m.

OpenStudy (anonymous):

ME means mechanical energy

OpenStudy (mony01):

Joules ?

OpenStudy (aum):

unit of energy. 1 Joule = 1 Newton-meter.

OpenStudy (mony01):

so 0.001 N-m

OpenStudy (aum):

Yes.

OpenStudy (mony01):

is that the answer?

OpenStudy (aum):

Yes.

OpenStudy (mony01):

thanks

OpenStudy (aum):

yw.

OpenStudy (perl):

nice work